# for Dynamic Neural Radiance Fields

### v2

@@ Abstract
The accurate reconstruction of dynamic scenes with neural radiance fields is significantly dependent on the estimation of camera poses. Widely used structure-from-motion pipelines encounter difficulties in accurately tracking the camera trajectory when faced with separate dynamics of the scene content and the camera movement. To address this challenge, we propose DynaMoN. DynaMoN utilizes semantic segmentation and generic motion masks to handle dynamic content for initial camera pose estimation and statics-focused ray sampling for fast and accurate novel-view synthesis. Our novel iterative learning scheme switches between training the NeRF and updating the pose parameters for an improved reconstruction and trajectory estimation quality. The proposed pipeline shows significant acceleration of the training process. We extensively evaluate our approach on two real-world dynamic datasets, the TUM RGB-D and the BONN RGB-D Dynamic dataset. DynaMoN improves over the state-of-the-art both in terms of reconstruction quality and trajectory accuracy. We plan to make our code public to enhance research in this area.
@@ I Introduction
To enable novel view synthesis using NeRF, accurate camera poses, usually retrieved via COLMAP [8] or the Apple ARKit, are essential. The structure-from-motion (SfM)-based approach often demands hours of computation to estimate reasonable camera poses and is challenged by large-scale dynamic scenes. To overcome the limitation of slow SfM approaches, recent works combined SLAM and NeRF [9, 10, 11, 2]. SLAM provides faster results for the camera trajectory than classic SfM. However, by nature, scenes are more often dynamic than static. Existing NeRF-SLAM [9, 10, 11, 2] methods assume a static scene which poses difficulties when dynamic scene content is present as shown in Fig 1.
Figure 1: NeRF-SLAM approaches rely on static scenes (left). Approaches such as InstantNGP [3] can be used to mask out dynamic scene content [2, 11] (left). Still, when using the masking, the scene representation lacks in quality (left). Fully dynamic NeRF provides an implicit 4D (3D+time) representation; however, it relies on offline SfM which can suffer in the presence of considerable motion (center). We consider dynamics via our motion-based segmentation (MS) and semantic segmentation (SS) already in the camera localization (right) and further during novel view synthesis. This enables a more robust camera tracking and novel view synthesis with a higher quality (ours, right).
In this paper, we present DynaMoN, a motion-aware camera localization and visualization approach which can handle highly dynamic scenes in camera pose retrieval and dynamic scene representation. We utilize semantic segmentation and generic motion masks during our camera localization process to enable robust tracking of the camera path in a dynamic environment. These masks are further utilized for ray sampling to train an efficient dynamic NeRF, based on HexPlane [4], that refines the initial camera poses, leading to improved reconstruction quality in scenes with strong motions.
@@ II-A Camera Localization and Scene Dynamics
In addition to classical or partially learnable SLAM approaches, Teed and Deng [16] introduce an end-to-end learnable method. Their approach, termed DROID-SLAM, consists of a dense bundle adjustment layer which enables recurrent iterative updates of camera poses and pixel-wise depth.
Typical camera localization approaches rely on a static scene assumption. However, natural scenes are often dynamic [18, 19, 20, 21]. To address this, Yu et al. [18] introduce a dynamic SLAM approach building upon ORB-SLAM2 [17]. Their DS-SLAM integrates a semantic segmentation network and a moving consistency check in the tracking process. Rünz et al. [22] enrich RGB-D SLAM with object-awareness and semantic segmentation capabilities while DynaSLAM [23] combines geometric and deep learning-based segmentation to mask dynamic objects in an attempt to ignore them for mapping. Others split the scene into dynamic and static parts. Dai et al. [20] build upon ORB-SLAM2 [17] and introduce a semantic segmentation using point correlation for this split and Ye et al. [21] divide the scene representation in a static flow field and a dynamic flow field.
@@ II-C Neural Representation of Dynamic Scenes
We, in contrast, split the problem into differentiation of static and dynamic content already for camera pose estimation. The dynamic environment is then represented by a dynamic, 4D NeRF calculated upon the camera poses from the static scene components, which is further utilized to refine these poses.
@@ III Method
We consider our input to be a video captured in a dynamic environment using a single monocular moving camera without pose knowledge. This leads to two objectives: first, estimating the camera pose, and second, reconstructing the scene which may comprise some deformation. We leverage masks to separate scene dynamics from static background for the former and utilize a 4D spatio-temporal neural representation for the latter. This allows to optimize for novel view synthesis while refining initial camera poses.
Figure 2: DynaMoN architecture. DynaMoN retrieves consecutive frames and applies semantic segmentation ($M_{ss}\rightarrow S_{S_{i}}$) and motion masks ($M_{MS}\rightarrow S_{M_{i}}$) on these input frames ($I$). Combining these masks (OR), we enable a motion-aware, fast and robust camera pose estimation. Based on the predicted camera poses $G$, we produce a time-dependent output ($\Theta$) using dynamic NeRF. During NeRF training, we utilize the previously computed masks to refine the camera poses (dashed lines). Dashed lines represent iterative updates on NeRF and pose parameters. We highlight our approach in green.
@@ III-A Robust Camera Localization in Dynamic Environments
Novel view synthesis methods commonly employ rigid SfM, often COLMAP [8], for camera pose estimation. This assumes a static scene. To cope with dynamics, it is possible to add binary masks to remove non-static regions. However, the high computation time remains a drawback. In our approach, we target both robustness against dynamics and a more efficient solution for pose estimation. Our underlying architecture is based on DROID-SLAM [16]. While it incorporates elements to cope with a certain level of scene dynamics, we found it can be improved by eliminating the contribution of the non-static regions for camera tracking. We do this by generating binary dynamics masks for each frame, which are combinations of semantic ($M_{SS}$) and motions masks ($M_{MS}$).
The semantic mask follows the paradigm that in real videos, the most prominent source of motion is often an object of known class. We therefore define $M_{SS}=\mathbb{1}(x>0)\circ\sum_{i}M_{SS}(i)$ for all defined semantic classes $i\in S$ of the set $S$ with the indicator function being $1$ for positive values. Additional moving objects are covered through a motion mask $M_{MS}$. One example is depicted by the moving box in Fig 2. Following DytanVO [34], our motion masks are generated using a CNN with input of previous image, forward optical flow and estimated camera motion to iteratively refine the mask along with pose updates. Instead of random initialization, we begin with the motion mask of the previous frame to improve stability and lower the number of iterations in our refinement. We combine both masks as $M=\mathbb{1}(x>0)\left(M_{SS}+M_{MS}\right)$.
III-B Neural Dynamic Scenes and Pose Refinement
Our scene representation builds on one of the recent dynamic NeRF models, HexPlane [4]. Targeting efficient optimization, fast rendering and high reconstruction quality, HexPlane [4] utilizes TensoRF [35] and constructs a 4D representation with three spatial and one temporal axis to create feature planes with their pairwise combinations to be the input of their tiny MLP. The grid size of each axis is governed in a coarse-to-fine fashion along the training to improve convergence and naturally regularizes the local neighborhoods. For optimization, it follows the original NeRF’s ray casting formulation and applies a pixel-wise photometric loss between the rendered RGB values $C$ and the ground truth $\hat{C}$, supported by a total variation regularization as:
@@ III-B Neural Dynamic Scenes and Pose Refinement
Sequences RGB-D Monocular
DVO-SLAM  [36] ORB-SLAM2 [17] PointCorr [20] DynaPix-D [37] DytanVO [34] DeFlowSLAM [21] DROID-SLAM [16] OURS (MS&SS) OURS (MS)
fr2/desk-person 0.104 0.006 0.008 - 1.166 0.013 0.006 0.007 0.006
fr3/sitting-static 0.012 0.008 0.010 - 0.016 0.007 0.005 0.005 0.005
fr3/sitting-xyz 0.242 0.010 0.009 - 0.260 0.015 0.009 0.010 0.009
fr3/sitting-rpy 0.176 0.025 0.023 - 0.046 0.027 0.022 0.024 0.021
fr3/sitting-halfsphere 0.220 0.025 0.024 - 0.310 0.025 0.014 0.023 0.019
fr3/walking-static 0.752 0.408 0.011 0.007 0.021 0.007 0.012 0.007 0.014
fr3/walking-xyz 1.383 0.722 0.087 0.014 0.028 0.018 0.016 0.014 0.014
fr3/walking-rpy 1.292 0.805 0.161 0.123 0.155 0.057 0.040 0.031 0.039
fr3/walking-halfsphere 1.014 0.723 0.035 0.023 0.385 0.420 0.022 0.019 0.020
Mean/Max 0.577/1.383 0.304/0.805 0.041/0.161 - 0.265/1.166 0.065/0.420 0.016/0.040 0.016/0.031 0.016/0.039
@@ III-C Pose Refinement and Training Strategy
We utilize eqn. (1) as the objective function for training both the NeRF and refining the camera poses. For the latter, we employ an alternating training strategy in which every 5,000 iterations the parameters to be optimized are switched between the radiance field and the pose parameters. To improve the accuracy and robustness of the optimization, during the pose refinement phase, we constrain the sampled rays, $\mathcal{R}$, to be elements of the static regions designated by the dynamics masks. Furthermore, we also use the same masks to ensure that the total variation regularization is imposed only on the static components. As the amount of non-dynamic pixels can vary from image to image, we accordingly scale the learning rate by the amount of static pixels in the image.
@@ IV Evaluation
DynaMoN is assessed on two challenging datasets, examining camera localization and novel view synthesis quality. To evaluate our approach, we consider two aspects. First, we consider the translational and rotational absolute trajectory error (ATE) [38] using root mean square error (RMSE) in meters and degrees. Second, we analyze the rendering quality of the novel dynamic views by reporting PSNR and SSIM [39] values.
TABLE II: RMSE of the translational ATE [m] on BONN RGB-D [13]. RGB-D results are reported by Palazzolo et al. [13]. For DROID-SLAM and ours, we calculated the results.
Sequences RGB-D Monocular
@@
To assess DynaMoN, we compare both camera localization and novel view synthesis part with state-of-the-art approaches on two challenging datasets. For the novel view synthesis results, we follow the usual splitting using every 8th frame of the TUM RGB-D [12] dataset and BONN RGB-D Dynamic [13] dataset for testing.
IV-A Implementation Details
We implement our approach in Python using PyTorch. For semantic masking, we use DeepLabV3 [41] with a ResNet50-backbone [42] trained on the large-scale COCO dataset [43] and use human masks. For motion masking, we use the trained model of DytanVO [34]. In practice, initialized with the motion mask of the previous frame, we found that setting the number of iterations for motion mask refinement to two provides stable convergence. For cases where the camera motion is higher than in the training sequences of DytanVO, the motion segmentation module results in an increase of false positives; as a counter-measure, we introduce a threshold for the number of pixels that contain motion. Consequently, masks exceeding this threshold are discarded. This ensures that there are always enough pixels for the dense bundle adjustment.
Regarding the loss function, we set $\lambda_{TV}$ initially to 0.005 and the number of sampled rays $\lvert\mathcal{R}\rvert$ per batch to 1024.
@@ IV-A Implementation Details
IV-B Camera Localization Quality
Our results on all datasets show a lower translational trajectory error compared to state-of-the-art approaches, see Table I and Table II. We compare RGB-D-based and monocular SLAM approaches with our proposed camera retrieval method using only motion masks (MS) as well as using motion and semantic segmentation masks in combination (MS&SS). Looking at the mean value of the translational trajectory error, our approach performs similarly to DROID-SLAM on TUM RGB-D. However, the maximum error value is much lower, which shows that it is less prone to failure when applied to real-world sequences. Moreover, comparing the rotational and translational ATE, see Figure LABEL:fig:atel, our MS&SS achieves lower values than DROID-SLAM. On the BONN RGB-D Dynamic dataset, our approach outperforms the state-of-the-art approaches in terms of mean and maximum trajectory error.
TABLE V: Analysis of the influence of different pose retrieval steps to PSNR and SSIM in the Bonn RGB-D Dynamics crowd2 scene with sparse input trajectory.
Approach DROID-SLAM + HexPlane
Our initial poses + HexPlane
DynaMoN
PSNR $\uparrow$
@@ IV-B Camera Localization Quality
16.46 / 0.661 17.07 / 0.660 17.31 / 0.659
IV-C Novel View Synthesis Quality
To evaluate our approach from a NeRF perspective we consider two aspects. First, the traditional way of generating camera ground truth when using NeRF, which is the use of COLMAP. Thus, we compare DynaMoN with HexPlane [4] using camera poses from COLMAP. Second, we highlight the influence of the iterative pose refinement by training both DynaMoN and HexPlane with our initially localized poses. Third, we compare our approach with RoDynRF [33] which regresses the camera pose along with a dynamic NeRF.

### v3

@@ Abstract
The accurate reconstruction of dynamic scenes with neural radiance fields is significantly dependent on the estimation of camera poses. Widely used structure-from-motion pipelines encounter difficulties in accurately tracking the camera trajectory when faced with separate dynamics of the scene content and the camera movement. To address this challenge, we propose Dynamic Motion-Aware Fast and Robust Camera Localization for Dynamic Neural Radiance Fields (DynaMoN). DynaMoN utilizes semantic segmentation and generic motion masks to handle dynamic content for initial camera pose estimation and statics-focused ray sampling for fast and accurate novel-view synthesis. Our novel iterative learning scheme switches between training the NeRF and updating the pose parameters for an improved reconstruction and trajectory estimation quality. The proposed pipeline shows significant acceleration of the training process. We extensively evaluate our approach on two real-world dynamic datasets, the TUM RGB-D dataset and the BONN RGB-D Dynamic dataset. DynaMoN improves over the state-of-the-art both in terms of reconstruction quality and trajectory accuracy. We plan to make our code public to enhance research in this area.
@@ I INTRODUCTION
To enable novel view synthesis using NeRF, accurate camera poses, usually retrieved via COLMAP [8] or the Apple ARKit, are essential. The structure-from-motion (SfM)-based approach often demands hours of computation to estimate reasonable camera poses and is challenged by large-scale dynamic scenes. To overcome the limitation of slow SfM approaches, recent works combined SLAM and NeRF [9, 10, 11, 2]. SLAM provides faster results for the camera trajectory than classic SfM. However, by nature, scenes are more often dynamic than static. Existing NeRF-SLAM [9, 10, 11, 2] methods assume a static scene which poses difficulties when dynamic scene content is present as shown in Fig 1.  [^1] This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
Figure 1: NeRF-SLAM approaches rely on static scenes (left). Approaches such as InstantNGP [3] can be used to mask out dynamic scene content [2, 11] (left). Still, when using the masking, the scene representation lacks quality (left). Fully dynamic NeRF provides an implicit 4D (3D+time) representation; however, it relies on offline SfM, which can suffer in the presence of considerable motion (center). We consider dynamics via our motion-based segmentation (MS) and semantic segmentation (SS) already in the camera localization (right) and further during novel view synthesis. This enables a more robust camera tracking and novel view synthesis with a higher quality (ours, right).
In this paper, we present DynaMoN, a motion-aware camera localization and visualization approach that can handle highly dynamic scenes in camera pose retrieval and dynamic scene representation. We utilize semantic segmentation and generic motion masks during our camera localization process to enable robust tracking of the camera path in a dynamic environment. These masks are further utilized for ray sampling to train an efficient dynamic NeRF, based on HexPlane [4], that refines the initial camera poses, leading to improved reconstruction quality in scenes with strong motions.
@@ II-A Camera Localization and Scene Dynamics
In addition to classical or partially learnable SLAM approaches, Teed and Deng [16] introduce an end-to-end learnable method. Their approach, termed DROID-SLAM, consists of a dense bundle adjustment layer that enables recurrent iterative updates of camera poses and pixel-wise depth.
Typical camera localization approaches rely on a static scene assumption. However, natural scenes are often dynamic [18, 19, 20, 21]. To address this, Yu et al. [18] introduce a dynamic SLAM approach building upon ORB-SLAM2 [17]. Their DS-SLAM integrates a semantic segmentation network and a moving consistency check in the tracking process. Rünz et al. [22] enrich RGB-D SLAM with object-awareness and semantic segmentation capabilities, while DynaSLAM [23] combines geometric and deep learning-based segmentation to mask dynamic objects in an attempt to ignore them for mapping. Others split the scene into dynamic and static parts. Dai et al. [20] build upon ORB-SLAM2 [17] and introduce a semantic segmentation using point correlation for this split and Ye et al. [21] divide the scene representation in a static flow field and a dynamic flow field. Yu et al. [24] combine semantic segmentation with optical flow to remove dynamic keypoints achieving a higher robustness compared to ORB-SLAM2.
@@ II-C Neural Representation of Dynamic Scenes
We, in contrast, split the problem into the differentiation of static and dynamic content already used for camera pose estimation. The dynamic environment is then represented by a dynamic, 4D NeRF calculated upon the camera poses from the static scene components, while further refining these poses based on the static parts of the images.
@@ III METHOD
We consider our input to be an RGB video captured in a dynamic environment using a single monocular moving camera without pose knowledge. This leads to two objectives: first, estimating the camera pose, and second, reconstructing the scene, which may comprise some deformation. We leverage masks to separate the scene dynamics from the static background for the former and utilize a 4D spatio-temporal neural representation for the latter. This allows to optimize for novel view synthesis while refining initial camera poses.
Figure 2: DynaMoN architecture. DynaMoN retrieves consecutive RGB frames and applies semantic segmentation ($M_{ss}\rightarrow S_{S_{i}}$) and motion masks ($M_{MS}\rightarrow S_{M_{i}}$) on these input frames ($I$). Combining these masks (OR), we enable a motion-aware, fast and robust camera pose estimation. Based on the predicted camera poses $G$, we produce a time-dependent output ($\Theta$) using dynamic NeRF. During NeRF training, we utilize the previously computed masks to refine the camera poses (dashed lines). Dashed lines represent iterative updates on NeRF and pose parameters. We highlight our approach in green.
@@ III-A Robust Camera Localization in Dynamic Environments
Novel view synthesis methods commonly employ rigid SfM, often COLMAP [8], for camera pose estimation. This assumes a static scene instead of dynamic scenes.
To cope with dynamics, it is possible to add binary masks to remove non-static regions. However, the high computation time of SfM approaches remains a drawback. In our approach, we target both robustness against dynamicscenes and a more efficient solution for pose estimation for dynamic moving cameras. Our underlying architecture is based on DROID-SLAM [16]. While it incorporates elements to cope with a certain level of scene dynamics, we found it can be improved by eliminating the contribution of the non-static regions for camera tracking. We do this by generating binary dynamics masks for each frame, which are combinations of semantic ($M_{SS}$) and motion masks ($M_{MS}$).
[…]
III-B Neural Dynamic Scenes
Our scene representation builds on a state-of-the-art dynamic NeRF, HexPlane [4]. Targeting efficient optimization, fast rendering and high reconstruction quality, HexPlane [4] utilizes TensoRF [36] and constructs a 4D representation with three spatial and one temporal axes to create feature planes with their pairwise combinations to be the input of their tiny MLP. The grid size of each axis is governed in a coarse-to-fine fashion along the training to improve convergence and naturally regularizes the local neighborhoods. For optimization, we follow the original NeRF ray casting and apply the pixel-wise photometric loss between the rendered RGB values $C$ and the ground truth values $\hat{C}$, supported by a total variation regularization as:
@@ III-B Neural Dynamic Scenes
Sequences RGB-D RGB
DVO-SLAM ORB-SLAM2 PointCorr DynaPix-D [37] Yu et al. DytanVO DeFlowSLAM DROID-SLAM OURS (MS&SS) OURS (MS)
[38] [17] [20] [37] [24] [35] [21] [16]
fr2/desk-person 0.104 0.006 0.008 - - 1.166 0.013 0.006 0.007 0.006
fr3/sitting-static 0.012 0.008 0.010 - 0.006 0.016 0.007 0.005 0.005 0.005
fr3/sitting-xyz 0.242 0.010 0.009 - - 0.260 0.015 0.009 0.010 0.009
fr3/sitting-rpy 0.176 0.025 0.023 - - 0.046 0.027 0.022 0.024 0.021
fr3/sitting-halfsphere 0.220 0.025 0.024 - 0.018 0.310 0.025 0.014 0.023 0.019
fr3/walking-static 0.752 0.408 0.011 0.007 0.009 0.021 0.007 0.012 0.007 0.014
fr3/walking-xyz 1.383 0.722 0.087 0.014 0.014 0.028 0.018 0.016 0.014 0.014
fr3/walking-rpy 1.292 0.805 0.161 0.123 0.031 0.155 0.057 0.040 0.031 0.039
fr3/walking-halfsphere 1.014 0.723 0.035 0.023 0.020 0.385 0.420 0.022 0.019 0.020
Mean/Max 0.577/1.383 0.304/0.805 0.041/0.161 - - 0.265/1.166 0.065/0.420 0.016/0.040 0.016/0.031 0.016/0.039
@@ III-C Pose Refinement and Training Strategy
We utilize eqn. (1) as the objective function for training both the NeRF and refining the camera poses. For the latter, we employ an alternating training strategy in which every 5 000 iterations, the parameters to be optimized are switched between the radiance field and the pose parameters. To improve the accuracy and robustness of the optimization, during the pose refinement phase, we constrain the sampled rays, $\mathcal{R}$, contributing to the loss to be elements of the static regions designated by the dynamics masks. Furthermore, we also use the same masks to ensure that the total variation regularization is imposed only on the static components. As the amount of non-dynamic sampled rays can vary from image to image, we scale the learning rate $\eta_{s}$ output by a learning rate scheduler by the fraction of static pixels in the image to keep the optimization steps in the same magnitude:
$\eta=\eta_{s}*\frac{n_{static}}{n_{static}+n_{dynamic}},$ (2)
where $n_{static}$ and $n_{dynamic}$ represent the number of static and dynamic pixels in the image and $\eta$ is the learning rate used for the pose refinement.
@@ IV EVALUATION
To evaluate our approach, we consider two aspects. First, we consider the translational and rotational absolute trajectory error (ATE) [39] using root mean square error (RMSE) in meters and degrees. Second, we analyze the rendering quality of the novel dynamic views by reporting PSNR and SSIM [40] values.
TABLE II: RMSE of the translational ATE [m] on BONN RGB-D Dynamic [13]. RGB-D results are reported by Palazzolo et al. [13]. For DROID-SLAM and ours, we calculated the results.
Sequences RGB-D RGB
@@
IV-A Datasets
To assess DynaMoN, we compare both the camera localization and the novel view synthesis part with state-of-the-art approaches on two challenging datasets.
TUM RGB-D - Dynamic subset [12]
The TUM RGB-D dataset has a resolution of $640\times 480$. While the dataset mainly focuses on static parts, five sequences were rated slightly dynamic and four highly dynamic.
BONN RGB-D Dynamic [13]
The BONN RGB-D Dynamic dataset consists of 24 dynamic sequences and two static sequences. The number of recordings per scene varies. The images have a resolution of $640\times 480$.
For the novel view synthesis results, we follow the usual splitting using every eighth frame of the TUM RGB-D [12] dataset and BONN RGB-D Dynamic [13] dataset for testing.
@@ IV-A Datasets
IV-B Implementation Details
We implemented our approach in Python using PyTorch and evaluated it on one single workstation (Intel i9-10980XE CPU, NVIDIA GeForce RTX 3090 GPU 24GB, and 128GB RAM) unless otherwise specified.
For semantic masking, we use DeepLabV3 [42] with a ResNet50-backbone [43] trained on the large-scale COCO dataset [44] and use human masks. For motion masking, we use the trained model of DytanVO [35]. In practice, initialized with the motion mask of the previous frame, we found that setting the number of iterations for motion mask refinement to two provides stable convergence. For cases where the camera motion is higher than in the training sequences of DytanVO, the motion segmentation module results in an increase of false positives; as a counter-measure, we introduce a threshold of 60 % for the maximum amount of pixels that contain motion. Consequently, masks exceeding this threshold are discarded to ensure that there are always enough pixels for the dense bundle adjustment. Additionally, we only consider a pixel as dynamic if the confidence of the network exceeds 0.95 for the first and 0.98 for the remaining mask refinement steps.
Regarding the loss function, we set $\lambda_{TV}$ initially to 0.005 and the number of sampled rays $\lvert\mathcal{R}\rvert$ per batch to 1 024. We employ the default optimization settings for the dynamic NeRF from [4] and use a multi-step learning rate decay scheduler for the pose refinement. The initial learning rate for the scheduler is set to $1*10^{-5}$ at the beginning of each pose optimization cycle and decayed every 1 000 iterations by a factor of $0.9$. As described beforehand, the actual used learning rate of the optimizer is retrieved by multiplication with the percentage of static pixels in the image. The number of optimization iterations is set to 200 000, unless otherwise stated.
TABLE V: Analysis of the influence of different pose retrieval steps to PSNR and SSIM in the Bonn RGB-D Dynamic crowd2 scene with sparse input trajectory.
Approach
@@ IV-B Implementation Details
DROID-SLAM [16] + HexPlane [4] 16.46 / 0.661
Our initial poses + HexPlane [4] 17.07 / 0.660
DynaMoN (Ours) 17.31 / 0.659
TABLE VI: Analysis of the runtime on the TUM RGB-D fr2/desk-person scene. $\ddagger$ means 100 000 training iterations and a batch size of 512.
$\dagger$ means 100 000 training iterations and a batch size of 1 024. The time format is hours:minutes:seconds.
[…]
RoDynRF [34] $\ddagger$ 01:52:55 08:00:44 00:11:15 10:04:54 6.05 / 0.127
DynaMoN (Ours) 00:13:49 09:49:25 00:42:32 10:45:46 24.83 / 0.725
DynaMoN (Ours) $\dagger$ 00:14:12 05:04:19 00:42:20 06:00:51 23.95 / 0.702
DynaMoN (Ours) $\ddagger$ 00:13:41 04:43:23 01:03:21 06:00:25 23.43 / 0.690
IV-C Camera Localization Quality
Our results on all datasets show a lower translational trajectory error compared to state-of-the-art approaches, see Table I and Table II. We compare RGB-D-based and monocular SLAM approaches with our proposed camera retrieval method using only motion masks (MS) as well as using motion and semantic segmentation masks in combination (MS&SS). Looking at the mean value of the RMSE of the translational trajectory error over all sequences, our approach performs similarly to DROID-SLAM on TUM RGB-D. However, the maximum error value is much lower, which shows that it is less prone to failure when applied to real-world scenarios. Moreover, comparing the RMSE of the rotational ATE, see Figure 3, our MS&SS-model achieves lower values than DROID-SLAM. On the BONN RGB-D Dynamic dataset, our approach outperforms the state-of-the-art approaches in terms of mean and maximum trajectory error for both rotation and translation.
IV-D Novel View Synthesis Quality
To evaluate our approach from a NeRF perspective we consider two aspects. First, the traditional way of generating camera ground truth when using NeRF is the use of COLMAP. Thus, we compare DynaMoN with HexPlane [4] using camera poses from COLMAP. Second, we highlight the influence of the iterative pose refinement by training both DynaMoN and HexPlane with our initially localized poses. Third, we compare our approach with RoDynRF [34] which regresses the camera pose along with a dynamic NeRF.